# Max-Rank: Efficient Multiple Testing for Conformal Prediction

### v3

@@ 6.1 Using max-rank for conformal prediction
Since we have intimately connected max-rank’s formulation to conformal prediction thus far, we may rightly assume that it integrates well with any CP algorithm following Algo. 1. In order to rigorously verify this fact, we assert that max-rank satisfies two key requirements imposed by conformal prediction: (1) the exchangeability condition on the underlying data is preserved, and (2) Eq. 7 is corrected in such a way that the conformal coverage guarantee from Eq. 2 is upheld.
We demonstrate in § A.4 that max-rank satisfies both conditions by looking at its two key operations, namely (1) operating in the rank space of nonconformity scores, and (2) applying a $\max$-operation to these ranks (or more generally, any set of exchangeable random variables). While the requirement on valid coverage directly follows from max-rank’s FWER-controlling property (Prop. 1), an additional proof is given via rank coverage. We summarize these insights as follows:
@@ 6.2 Conformal prediction experiments
Finally, we showcase max-rank’s usefulness experimentally for two applications, conformal multi-target regression and conformal object detection.
Experiment design. We consider applying conformal prediction to multi-target regression problems where $m$ real-valued and possibly dependent targets are predicted simultaneously. We aim to ensure that all $m$ targets are jointly covered with probability $(1-\alpha)$ by our conformal procedure, and consider a range of multiple testing corrections for that purpose: Bonferroni and its slightly more powerful Šidák-adjustment (Abdi et al., 2007), two copula-based corrections from Messoudi et al. (2021), max-T (Westfall and Young, 1993) and finally max-rank. We include the copula-based corrections because they also address settings of positive dependency, and since max-rank’s theoretical motivation via copulas provides an interesting link. For CP (Algo. 1) we employ simple nonconformity scores of the form $|Y_{i,k}-\hat{f}(X_{i,k})|$ where $i\in[n],k\in[m]$, and construct the prediction intervals for test points as $\hat{C}(X_{n+1,k})=[\hat{f}(X_{n+1,k})-\hat{Q}_{1-\hat{\alpha},k},\,\hat{f}(X_{n+1,k})+\hat{Q}_{1-\hat{\alpha},k}]$, where $\hat{Q}_{1-\hat{\alpha},k}$ now denotes the conformal quantile at a possibly adjusted individual significance level $\hat{\alpha}_{k}$. We compare results with different corrections across two key conformal metrics (Vovk et al., 2005): empirical coverage as a measure of validity to verify Eq. 2, and average prediction interval width to assess the efficiency of the solution. See also Appendix D.
@@
Table 1: Empirical coverage and mean prediction interval width for different corrections ($\alpha=0.1$) on three multi-target datasets: scpf (1137 samples), rf1 (9125 samples), and scm1d (9803 samples). We employ a random forest regressor (Pedregosa et al., 2011) and $20\%$ calibration split, and report means and std. deviation across $100$ random trials. max-rank consistently provides strong results (best results, second best).
Conformal multi-target regression. Following Messoudi et al. (2021) we consider three tabular datasets with a varying number of targets [^4] Obtainable at https://github.com/tsoumakas/mulan: scpf (website traffic data with $m=3$), rf1 (river flow data with $m=8$), and scm1d (product pricing data with $m=16$). Our results for the usual target coverage of $(1-\alpha)=90\%$ are displayed in Table 1. Firstly, we observe that all corrections provide valid coverage, whereas without we (strongly) undercover. Among corrections, Bonferroni and Bonferroni-Šidák provide overly conservative results, and copula methods perform generally well. In particular for the scpf dataset the parametric Gumbel copula benefits from the low test dimensionality ($m=3$), while max-T and max-rank suffer from its small sample size. Overall, max-rank tends to outperform other corrections in terms of interval width, and matches the target coverage level very closely.
@@ 7 DISCUSSION
We introduced max-rank, an efficient MHT correction for testing under positive dependency which tailors Westfall and Young (1993)’s corrections for conformal prediction. While we provide strong theoretical and empirical results, our method’s practicality is restricted to (1) data dependencies that become apparent through rank ordering, (2) settings of positive dependency, since we do not benefit from rank information under negative dependency and thus do not improve upon Bonferroni, and (3) its permutation-based, data-driven nature necessitating a sufficient ratio of test statistics $n$ to test dimensions $m$ for empirical FWER control. Future work includes establishing stronger theoretical results, such as on convergence rates to target level $\alpha$, the dependence between quantities $m$, $n$, and $\alpha$, and its relationship to the empirical copula. Of course, we may also experimentally extend to further applications such as multi-step time series (Sun and Yu, 2023). Our work contributes to the ongoing research on multiple testing under dependency, all the while addressing existing needs in predictive settings using conformal prediction.
@@ Acknowledgements
We thank Aaditya Ramdas and Jelle Goeman for making us aware of connections to the Westfall & Young testing procedures. This project was generously supported by the Bosch Center for Artificial Intelligence. Eric Nalisnick did not use any resources from Johns Hopkins for this work. Author contributions are as follows: Alexander Timans established the theoretical connections to conformal prediction and Westfall & Young, performed experiments and paper writing. Christoph-Nikolas Straehle developed the max-rank algorithm and proved several theoretical properties. Kaspar Sakmann, Christian Naesseth and Eric Nalisnick provided project guidance as well as technical and writing feedback.
@@ Appendix A MATHEMATICAL DETAILS
This section includes the details of our theoretical results for max-rank (Prop. 1 and Prop. 2 in § 4), as well as its use for conformal prediction (Prop. 3 in § 6.1).
@@ A.2 Obtaining a solution via max-rank
For a given solution $q_{*}$ satisfying Eq. 10, the desired quantile vector ensuring FWER control is then computed by evaluating the probability under each inverse marginal as $\mathbf{x_{*}}=(F_{X_{1}}^{-1}(q_{*}),\dots,F_{X_{m}}^{-1}(q_{*}))\in\mathbb{R}^{m}$. Note that now the elements of $\mathbf{x_{*}}$ may differ from each other, since the marginals are not necessarily identical. The obtained quantiles in $\mathbf{x_{*}}$ are equal to the adjusted quantiles $\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m}$ described in Algo. 2. By relating the element-wise ranks of $x_{*,k}$ under $\mathbf{S}$ to the marginals $F_{X_{k}}$ and establishing a correspondence between the copula and a joint c.d.f. in rank space, we next show that our max-rank correction provides a solution $q_{\max}$ to Eq. 10.
For that, we relate the algorithmic steps in Algo. 2 to our problem setting in § A.1. Consider the matrix of test statistics $\mathbf{S}\in\mathbb{R}^{n\times m}$ first introduced in § 3, and denote by $S_{k}$ its $k$-th column corresponding to the $k$-th empirical null. We then associate $S_{k}$ with the $k$-th dimension of $\mathbf{X}$, i.e., the elements $s_{i,k}\in S_{k},\,i\in[n]$ match to $n$ realizations $\{x_{1,k},\dots,x_{n,k}\}$ of $X_{k}$, and $S_{k}$ thus provides an empirical estimate of the marginal $F_{X_{k}}$. Next, we evaluate the element-wise ranks of $\mathbf{x_{*}}$ under $\mathbf{S}$. For any element $x_{*,k}\in\mathbf{x_{*}}$ a direct correspondence between its rank under $S_{k}$ (see Eq. 5) and marginal $F_{X_{k}}$ is given by
@@
where $F_{\max\mathbf{\tilde{R}}}$ now denotes the one-dimensional c.d.f. in the space of maximum ranks across dimensions $k\in[m]$. Intuitively, ensuring that the condition holds jointly for ranks in all dimensions is equivalent to satisfying it in the ‘worst’ one. $F_{\max\mathbf{\tilde{R}}}$ is precisely the space in which max-rank operates after the $l^{\infty}$-norm is applied, and in which following Algo. 2 a rank $\tilde{r}_{\max}=r_{\max}/n=q_{\max}$ with probability mass $(1-\alpha)$ below it is computed. Thus we can employ Eq. 12 in its reverse direction to claim
@@
therefore providing a solution which satisfies the constraint in Eq. 10. Furthermore, since we precisely compute the rank at target level $(1-\alpha)$ (and ignoring any finite-sample correction), our solution $q_{\max}$ also satisfies the $\min$ in Eq. 10, and the obtained quantiles $\mathbf{x}_{\max}$ are optimally tight in terms of coverage under the imposed restriction that $q_{\max}=q_{\max,1}=\dots=q_{\max,m}$. Note that if the restriction is lifted, marginally tighter individual quantiles $x_{\max,k}$ are potentially obtainable in some dimensions. However, it becomes unclear how the required algorithmic modifications to max-rank would continue to uphold FWER control.
@@ A.4 Using max-rank for conformal prediction
Two key conditions are imposed on any multiple testing correction that integrates with conformal prediction: (1) the exchangeability condition on the underlying data is preserved, and (2) Eq. 7 is corrected in such a way that the conformal coverage guarantee is upheld. We next show that max-rank satisfies both conditions by looking at its two key operations, namely (1) operating in the rank space of nonconformity scores, and (2) applying a $\max$-operation to these ranks. We first address the exchangeability requirement, followed by validity.
@@
For the $\max$-operation, we make a simple argument on stochastic dominance of the quantiles. By definition of the composite empirical null $\mathbf{r}_{\max}$ in Algo. 2 we have for any $k\in[m]$ that $\hat{Q}(1-\alpha;\mathbf{r}_{\max})\geq\hat{Q}(1-\alpha;\mathbf{r}_{k})$ holds. Clearly, a prediction set $\hat{C}_{k}$ constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{k})$ is then a subset of $\hat{C}_{\max}$ constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$, i.e., $\hat{C}^{\max}\supseteq\hat{C}^{k}$. Since by Eq. 4 and validity of operating in the rank space $\hat{C}_{k}$ is constructed to ensure target coverage, the same coverage guarantee applies to $\hat{C}_{\max}$.

### v4

@@ 6.1 Using max-rank for conformal prediction
Since we have closely connected max-rank’s formulation to conformal prediction, one can rightly assume that it integrates well with any conformal procedure following Algo. 1. To rigorously assert this, we verify that max-rank meets two key requirements of conformal prediction: (i) the exchangeability condition on the underlying data is preserved, and (ii) Eq. 7 is corrected in such a way that the conformal coverage guarantee from Eq. 2 is upheld.
We demonstrate in § A.4 that max-rank satisfies both conditions by analysing its two key operations, namely (i) operating in the rank space of nonconformity scores, and (ii) applying a $\max$-operation on these ranks (or more generally, any set of exchangeable random variables). While the requirement on valid coverage technically directly follows from max-rank’s FWER-controlling property (Prop. 1), an additional proof is provided via rank coverage. We summarize these insights in the following proposition:
@@ 6.2 Conformal prediction experiments
Table 1: Empirical coverage and mean prediction interval width for different corrections ($\alpha=0.1$) on three multi-target datasets: scpf (1137 samples), rf1 (9125 samples), and scm1d (9803 samples). We employ a random forest regressor (Pedregosa et al., 2011) and $20\%$ calibration split, and report means and std. deviation across $100$ random trials. max-rank consistently provides strong results (best, second best results).
Finally, we showcase max-rank’s usefulness experimentally for two applications, conformal multi-target regression and conformal object detection. Our code is available at https://github.com/alextimans/max-rank.
Experiment design. We consider applying conformal prediction to multi-target regression problems where $m$ real-valued and possibly dependent targets are predicted simultaneously. We aim to ensure that all $m$ targets are jointly covered with probability $(1-\alpha)$ by our conformal procedure, and consider a range of multiple testing corrections for that purpose: Bonferroni and its slightly more powerful Šidák-adjustment (Abdi et al., 2007), two copula-based corrections from Messoudi et al. (2021), max-T (Westfall and Young, 1993) and finally max-rank. We include the copula-based corrections because they also address settings of positive dependency, and since max-rank’s theoretical motivation via copulas provides an interesting link. For CP (Algo. 1) we employ nonconformity scores of the simple residual form $s_{i,k}=|Y_{i,k}-\hat{f}(X_{i,k})|$ where $i\in[n],k\in[m]$, and construct the prediction intervals for test points as $\hat{C}(X_{n+1,k})=[\hat{f}(X_{n+1,k})-\hat{Q}_{1-\hat{\alpha}_{k}},\,\hat{f}(X_{n+1,k})+\hat{Q}_{1-\hat{\alpha}_{k}}]$, where $\hat{Q}_{1-\hat{\alpha}_{k}}$ again denotes the conformal quantile at adjusted individual significance level $\hat{\alpha}_{k}$ using a MHT correction. We compare results with different corrections across two key conformal metrics (Vovk et al., 2005): empirical coverage as a measure of validity to verify Eq. 2, and average prediction interval width to assess the efficiency of the solution. See also Appendix D.
Conformal multi-target regression. Following Messoudi et al. (2021) we consider three tabular datasets with a varying number of targets [^3] Obtainable at https://github.com/tsoumakas/mulan: scpf (website traffic data with $m=3$), rf1 (river flow data with $m=8$), and scm1d (product pricing data with $m=16$). Our results for the usual target coverage of $(1-\alpha)=90\%$ are displayed in Table 1. Firstly, we observe that all corrections provide valid coverage, whereas without we strongly undercover. Among corrections, Bonferroni and Bonferroni-Šidák provide overly conservative results, and copula methods perform generally well. In particular for the scpf dataset the parametric Gumbel copula benefits from the low test dimensionality ($m=3$), while max-T and max-rank suffer from its small sample size. Overall, max-rank tends to outperform other corrections in terms of interval width, and matches the target coverage level very closely.
@@ 7 DISCUSSION
We introduce max-rank, an efficient MHT correction for testing under positive dependency which adapts Westfall and Young (1993)’s corrections to conformal prediction. While we provide strong theoretical and empirical results, our method’s benefits are mostly restricted to (i) data dependencies that become apparent through rank ordering, (ii) settings of positive dependency, since we do not benefit from rank information under negative dependency and thus do not improve upon Bonferroni, and (iii) its permutation-based, data-driven nature necessitates a sufficient ratio of test statistics $n$ to test dimensions $m$ for empirical FWER control. Future work includes establishing stronger theoretical results, such as on convergence rates to target level $\alpha$, the dependence between quantities $m$, $n$, and $\alpha$, and its relationship to the empirical copula. Naturally, we may also experimentally extend to further applications such as multi-step time series (Sun and Yu, 2023). We hope that max-rank contributes to the ongoing research on multiple testing under dependency, all the while addressing practical needs in predictive settings using conformal prediction.
@@ Acknowledgements
We thank Aaditya Ramdas and Jelle Goeman for making us aware of connections to the Westfall & Young testing procedures, and Drew Prinster for helpful feedback on an earlier draft. This project was generously supported by the Bosch Center for Artificial Intelligence. Eric Nalisnick did not use any resources from Johns Hopkins for this work. Author contributions are as follows: Alexander Timans established the theoretical connections to conformal prediction and Westfall & Young, performed experiments and paper writing. Christoph-Nikolas Straehle developed the max-rank algorithm and proved several theoretical properties. Kaspar Sakmann, Christian A. Naesseth and Eric Nalisnick provided project guidance and helpful feedback.
CHECKLIST
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] Algorithmic descriptions are included in the paper, the full proofs are included in the appendix.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] We discuss the algorithmic properties and empirically verify sample size and runtime requirements, but do not include a formal theoretical analysis.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] Our code is publicly available at https://github.com/alextimans/max-rank.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] The full proofs are included in the appendix.
    2. (b)
      Complete proofs of all theoretical results. [Yes] The full proofs are included in the appendix.
    3. (c)
      Clear explanations of any assumptions. [Yes] The full proofs are included in the appendix.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] We provide experimental design details in the paper and the appendix, and the code is publicly available.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] All results are reproducible on a single consumer-grade GPU (e.g., NVIDIA RTX 4090). max-rank is computationally light-weight and is easily runnable on CPU for smaller experiments.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes] Our code is publicly released under an AGPL-3.0 license.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes] The code is available via an URL.
    4. (d)
      Information about consent from data providers/curators. [Not Applicable] Used datasets are synthetic or publicly available.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Max-Rank: Efficient Multiple Testing for Conformal Prediction
– Supplementary Materials –
The appendix is organized as follows:
  • Mathematical details and proofs in Appendix A
  • Algorithmic details on max-T and min-P in Appendix B
  • Additional background on FDR control and the Bonferroni correction in Appendix C
  • Additional experiment details for our simulation study, multiple testing baselines, metrics, and conformal experiments in Appendix D
Notation Type Description
$n$ $\in\mathbb{N}^{+}$ Nr. of calibration samples, nr. of test statistics, index $i=1,\dots,n$
$m$ $\in\mathbb{N}^{+}$ Nr. of conformal procedures, nr. of test dimensions, index $k=1,\dots,m$
$\alpha,\alpha_{k}$ $\in(0,1)$ Target miscoverage rate, target significance level (global, in the $k$-th dimension)
$\hat{\alpha},\hat{\alpha}_{k}$ $\in(0,1)$ Adjusted or corrected significance level (global, in the $k$-th dimension)
$\hat{f}(\cdot)$ $\mathcal{X}\rightarrow\mathcal{Y}$ Pre-trained prediction model
$s(\cdot)$ $\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ Conformal scoring function
$s_{i},s_{i,k}$ $\in\mathbb{R}$ Nonconformity score for the $i$-th sample, test statistic (in the $k$-th dimension)
$r_{i},r_{i,k}$ $\in\mathbb{N}$ Rank of nonconformity score for the $i$-th sample (in the $k$-th dimension)
$S$ $\subset\mathbb{R}$ Set of scores $\{s_{1},\dots,s_{n}\}$ given a single dimension (equiv. to fixed $\mathbf{s}_{k}$ ignoring ordering)
$\hat{C}(X_{n+1})$ $\subseteq\mathcal{Y}$ Conformal prediction set for test input $X_{n+1}$
$\hat{P}_{n+1}(y;S)$ $\in[0,1]$ Conformal p-value for the null hypothesis $H_{0}:Y_{n+1}=y$ based on $S$
$\hat{Q}(1-\alpha;S),\,\hat{Q}_{1-\alpha}$ $\in\mathbb{R}$ Conformal quantile at target coverage $(1-\alpha)$ based on $S$
$\hat{Q}(1-\alpha_{k};\mathbf{s}_{k}),\hat{Q}_{1-\alpha_{k}}$ $\in\mathbb{R}$ Conformal quantile at target coverage $(1-\alpha_{k})$ based on $\mathbf{s}_{k}$ (i.e., in the $k$-th dimension)
$\mathbf{s}_{k}$ $\in\mathbb{R}^{n}$ Vector of scores, test statistics, empirical null distribution (in the $k$-th dimension)
$\mathbf{r}_{k}$ $\in\mathbb{N}^{n}$ Vector of score ranks, empirical null distribution in rank space (in the $k$-th dimension)
$\mathbf{S}$ $\in\mathbb{R}^{n\times m}$ Matrix of test statistics with columns $\mathbf{s}_{1},\dots,\mathbf{s}_{m}$
$\mathbf{R}$ $\in\mathbb{N}^{n\times m}$ Matrix of column-wise ranks of $\mathbf{S}$
$\mathbf{r}_{\max}$ $\in\mathbb{N}^{n}$ Vector of max. score ranks, composite empirical null distribution in rank space
$r_{\max}$ $\in\mathbb{N}$ The rank value $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ representing the corrected significance levels $\hat{\alpha},\hat{\alpha}_{k}$
Table 2: Notation table with clarifications for key quantities presented in the main paper.
@@ Appendix A MATHEMATICAL DETAILS
This section includes notation clarification in Table 2, and the details of our theoretical results for max-rank (Prop. 1 and Prop. 2 in § 4), as well as its use for conformal prediction (Prop. 3 in § 6.1).
@@ A.2 Obtaining a solution via max-rank
For a given solution $q_{*}$ satisfying Eq. 10, the desired quantile vector ensuring FWER control is then computed by evaluating the probability under each inverse marginal as $\mathbf{x_{*}}=(F_{X_{1}}^{-1}(q_{*}),\dots,F_{X_{m}}^{-1}(q_{*}))\in\mathbb{R}^{m}$. Note that now the elements of $\mathbf{x_{*}}$ may differ from each other, since the marginals are not necessarily identical. The obtained quantiles in $\mathbf{x_{*}}$ are equal to the adjusted quantiles $\hat{Q}_{1-\hat{\alpha}_{1}},\dots,\hat{Q}_{1-\hat{\alpha}_{m}}$ described in Algo. 2. By relating the element-wise ranks of $x_{*,k}$ under $\mathbf{S}$ to the marginals $F_{X_{k}}$ and establishing a correspondence between the copula and a joint c.d.f. in rank space, we next show that our max-rank correction provides a solution $q_{\max}$ to Eq. 10.
For that, we relate the algorithmic steps in Algo. 2 to our problem setting in § A.1. Consider the matrix of test statistics $\mathbf{S}\in\mathbb{R}^{n\times m}$ first introduced in § 3, and denote by $S_{k}$ its $k$-th column corresponding to the $k$-th empirical null (in set notation, rather than as a vector $\mathbf{s}_{k}$). We then associate $S_{k}$ with the $k$-th dimension of $\mathbf{X}$, i.e., the elements $s_{i,k}\in S_{k},\,i\in[n]$ match with the set of $n$ realizations $\{x_{1,k},\dots,x_{n,k}\}$ of $X_{k}$, and $S_{k}$ thus provides an empirical estimate of the marginal $F_{X_{k}}$. Next, we evaluate the element-wise ranks of $\mathbf{x_{*}}$ under $\mathbf{S}$. For any element $x_{*,k}\in\mathbf{x_{*}}$ a direct correspondence between its rank under $S_{k}$ (see Eq. 5) and marginal $F_{X_{k}}$ is given by
@@
where $F_{\max\mathbf{\tilde{R}}}$ now denotes the one-dimensional c.d.f. in the space of maximum ranks across dimensions $k\in[m]$. Intuitively, ensuring that the condition holds jointly for ranks in all dimensions is equivalent to satisfying it in the ‘worst’ one. $F_{\max\mathbf{\tilde{R}}}$ is precisely the space in which max-rank operates after the $l^{\infty}$-norm is applied (matching the composite null $\mathbf{r}_{\max}$ in the paper), and in which following Algo. 2 a rank $\tilde{r}_{\max}=r_{\max}/n=q_{\max}$ with probability mass $(1-\alpha)$ below it is computed. Thus we can employ Eq. 12 in its reverse direction to claim
@@
therefore providing a solution which satisfies the constraint in Eq. 10. Furthermore, since we precisely compute the rank at target level $(1-\alpha)$ (and ignoring any finite-sample correction), our solution $q_{\max}$ also satisfies the $\min$ in Eq. 10, and the obtained quantiles $\mathbf{x}_{\max}$ are optimally tight in terms of coverage under the imposed restriction that $q_{\max}=q_{\max,1}=\dots=q_{\max,m}$. Note that if the restriction is lifted, marginally tighter individual quantiles $x_{\max,k}$ are potentially obtainable in some dimensions. However, it becomes unclear how the required algorithmic modifications to max-rank would continue to formally uphold FWER control.
@@ A.4 Using max-rank for conformal prediction
Two key conditions are imposed on any multiple testing correction that integrates with conformal prediction: (i) the exchangeability condition on the underlying data is preserved, and (ii) Eq. 7 is corrected in such a way that the conformal coverage guarantee is upheld. We next show that max-rank satisfies both conditions by looking at its two key operations, namely (i) operating in the rank space of nonconformity scores, and (ii) applying a $\max$-operation to these ranks. We first address the exchangeability requirement, followed by validity.
@@
For the $\max$-operation, we make a simple argument on stochastic dominance of the quantiles. By definition of the composite empirical null $\mathbf{r}_{\max}$ in Algo. 2 we have for any $k\in[m]$ that $\hat{Q}(1-\alpha;\mathbf{r}_{\max})\geq\hat{Q}(1-\alpha;\mathbf{r}_{k})$ holds. Clearly, a prediction set $\hat{C}_{k}$ for the $k$-th dimension constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{k})$ is then a subset of the prediction set $\hat{C}_{\max}$ in the $k$-th dimension constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$, i.e., $\hat{C}_{\max}\supseteq\hat{C}_{k}$. Since by Eq. 4 and validity of operating in the rank space $\hat{C}_{k}$ is constructed to ensure target coverage, the same coverage guarantee applies to $\hat{C}_{\max}$.